Any help with this one will be appreciated!
What is the coefficient of x^{46} in the expansion of(2x-1)^{100}?

Answer by rothauserc(4718)

coefficient of x^46 in expansion of (2x-1)^100
:
we use the binomial expansion theorem
:
a=2x, b=-1, 100-46=54
:
coefficient of x^46 = 100C54 * (2x)^(100-54) * (-1)^54 =
:
73.470998191e+27 * 70.368744178e+12 * 1 =
:
***************
5.170061876e+42
***************
: